OpenStudy (anonymous):

Find the resistance of a circular coil of 500 turns, the radius of each of them is 7.5 cm. If the radius of the wire itself is 0.01mm, its Resistivity is 2x10^-8 ohm.m.

OpenStudy (shamim):

u hv to use formula R=rho*L/A

OpenStudy (anonymous):

r = radius, L = length, A = area? I only have the radius of each turn, the radius of the wire and resistivity. What is h and o? Can yo please explain more?

OpenStudy (shamim):

i mean rho=resistivity

OpenStudy (anonymous):

Aha, how can I get L and A then?

OpenStudy (shamim):

ok

OpenStudy (shamim):

wts the circumference of a circlr of wire

OpenStudy (anonymous):

The radius of each turn is 7.5cm, there are 500 turns. Also the radius of the wire is 0.01mm, I'm not sure what to do with the radius of the wire. But I'm guessing I can get the circumference with \[2*\pi*r\]

OpenStudy (shamim):

ya u r right here r=7.5cm

OpenStudy (shamim):

u r going to find out the circumference of circle

OpenStudy (shamim):

n radius of that circle is r=7.5cm

OpenStudy (shamim):

bt need to convert the unit frm cm to meter

OpenStudy (shamim):

right?

OpenStudy (anonymous):

right and then?

OpenStudy (shamim):

can u find out the cross sectional area of ur given wire

OpenStudy (shamim):

i mean A=?

OpenStudy (anonymous):

My problem is that I don't know what to do with 2 radii. To find A I guess we will use \[\pi*r ^{2}\]

OpenStudy (shamim):

ya ur guess is almost correct

OpenStudy (shamim):

for calculating cross sectional area A=π*r^2 here u hv to use r=0.01mm

OpenStudy (shamim):

because 0.01mm is the radius of wire not circle

OpenStudy (shamim):

right?

OpenStudy (shamim):

but u hv to convert the unit frm mm to meter

OpenStudy (anonymous):

I'm not sure when to use the radius 7.5cm and when to use 0.01mm

OpenStudy (shamim):

when u r going to calculate the circumference of ur given circle of wire, u hv to use the radius of that circle n its bigger r=7.5cm

OpenStudy (shamim):

anyway if u feel confused then we can talk over skype, if u want

OpenStudy (anonymous):

Thanks to you I managed to solve it!